USA Today article claims that the proportion of people who believe global warming is a serious issue is 0.52, but given the number of people you've talked to about this same issue, you believe it is less than 0.52. The hypotheses for this test are Null Hypothesis: p ≥ 0.52, Alternative Hypothesis: p < 0.52. If you randomly sample 29 people and 17 of them believe that global warming is a serious issue, what is your test statistic and p-value?

Question 11 options:

1) Test Statistic: 0.714, P-Value: 1.524

2) Test Statistic: -0.714, P-Value: 0.238

3) Test Statistic: -0.714, P-Value: 0.762

4) Test Statistic: 0.714, P-Value: 0.238

5) Test Statistic: 0.714, P-Value: 0.762
Answer #1

Sol:

Null Hypothesis: p ≥ 0.52

Alternative Hypothesis: p < 0.52

test statistic

z=p^-p/sqrt(p*(1-p)/n

p^=x/n=17/29=0.5862069

n=29

z=(0.5862069-0.52)/sqrt(0.52*(1-0.52)/29)

z= 0.7136413

z=0.714

pvalue in excel

=NORM.S.DIST(0.714;TRUE)

=0.76238641

p=0.762

MARK OPTION 5